OpenStudy (anonymous):

write the equation of the line with the given y-intercept and passing through the given point. Leave answer in slope intercept form. (0,-3) (4,-6) -6+ (-3) = -3/4 4-0 y=-3/4 -3 is this correct

OpenStudy (anonymous):

yes: but y = -3/4 x - 3

OpenStudy (anonymous):

write the equation of a line in slope intercept form with the given slope and passing through the given point. m =-4 (-3,0) y -0=-4(x+3) y=-4x-12 is this correct

OpenStudy (anonymous):

correct